OpenStudy (anonymous):

Simplify the following expression by adding and subtracting: 4√12 + 7√48 - 2√8 A. -8√2 + 228√3 B. -6√2 + 31√3 C. -4√2 + 17√3 D. -4√2 + 36√3

OpenStudy (paki):

i solved this but my answer is something else... =4√12+7√48-2√8 =4*2√3+7*4√3-2*2√2 =8√3+28√3-4√2 = - 4√2 + 36√3 ...... answer.... option D is rite one

OpenStudy (paki):

@Abiilovee911